OpenStudy (sparklestaraa):

Jose is comparing the prices of two carpet cleaning companies. Company A charges $0.40 per square foot of carpet area and an additional $22 per day for services. Company B charges 0.60 per square foot of carpet area and an additional $18 per day for services. Part A: Write equations to represent Company A's and Company B's total charges for cleaning a certain carpet area in a day. Define the variable used in the equations. Part B: Which company would charge less for cleaning 500 square feet of carpet area? Justify your answer. Part C: How much money is saved by using

OpenStudy (sparklestaraa):

@dtan5457

OpenStudy (godlovesme):

company a chages 0.40 per + 22 --> y=0.40x +22 make sense

OpenStudy (dtan5457):

Make 2 equations here

OpenStudy (dtan5457):

Company A: y=0.40x+22 COMPANY B: y=0.60x+18

OpenStudy (dtan5457):

x=the amount of square feet of carpet

OpenStudy (dtan5457):

"Part B: Which company would charge less for cleaning 500 square feet of carpet area? Justify your answer." Plug 500 into each x value of each equation and find your output (y)

OpenStudy (dtan5457):

Can you do that for me?

OpenStudy (sparklestaraa):

yes hold on

OpenStudy (dtan5457):

Got it yet?

OpenStudy (sparklestaraa):

Yes i do :D its y=318 @dtan5457

OpenStudy (sparklestaraa):

but which company?

OpenStudy (sparklestaraa):

this is overall what i wrote for the answer- Part A- Company A: y=0.40x +22 Company B: y=0.60x+18 x=the amount of square feet of carpet. Part B- Company A: y=0.40x +22 y=0.40(500)+22 y=222 Company B: y=0.60x+18 y=0.60(500)+18 y=318 Company A would charge less because $222 dollars is less than $318 dollars for company B's services. Part C- Company A= $122 Company B= $138 $138-$122= $16 $16 dollars is saved.

OpenStudy (dtan5457):

318-222=???

OpenStudy (dtan5457):

For part c

OpenStudy (sparklestaraa):

200 sq ft for x in the last one right?

OpenStudy (dtan5457):

part c is much money is saved. company a is $222 company b is $318 $318-$222=how much money saved

OpenStudy (sparklestaraa):

ohhhh see because i read it as How much money is saved by using the services of Company A instead of Company B for cleaning 200 square feet (not 500) <<<<<< of carpet area?

OpenStudy (dtan5457):

yes its 500

OpenStudy (sparklestaraa):

ohh ok :D thankss